# Centre du Parc, Rue Marconi 19, CH-1920 Martigny, Switzerland

### v1

@@ 6 Discussion
The primary objective of this article is to enhance the sampling distribution employed in SMPC through the application of products of experts. As anticipated, both methodologies (NFs-PoE-MPPI and TT-PoE-MPPI) exhibit superior performance when compared to the conventional MPPI method. However, the performance of NFs-PoE-MPPI was not significantly better than that of Proj-MPPI, and in all experiments, TT-PoE-MPPI consistently outperformed NFs-PoE-MPPI by a substantial margin. In the subsequent sections, we will delve deeper into the reasons behind this performance, explaining the advantages of using TT-distribution over NFs models and identifying its limiting factors.
@@ 6.1 TT-distribution vs Normalizing flow
Nevertheless, similar to other latent space learning methods, the acquired latent space lacks intuitiveness (Losey et al., 2022). In this context, intuitive refers to the observed relation between data points in the latent space, not necessarily aligning with their distance in the main spaces. For instance, when two proximate data points exhibit similar distances to an objective point in latent space, while in task space, one is positioned near the nominal solution and the other is distant. This phenomenon is illustrated in Fig. 11, obtained from a model trained for the PNGRID task, where the nominal action is presumed to be $[-0.5,-0.5]$ at state $[-1,-0.9]$, and the Gaussian model utilized for the MPPI method possesses identical variance in both directions. At this point, only velocity constraints (ranging between -1 and 1) must be active, with no other constraints in effect.
@@
Observing the task space samples, it is evident that they are not entirely clustered around the nominal action $[-0.5,0.5]$. Additionally, the distribution of samples in both directions appears dissimilar. This discrepancy arises due to the nonlinear function employed to revert the data back to the task space, which may have varying effects in different directions. This suggests that the covariance matrix in the latent space must be meticulously chosen to ensure the desired distribution upon reverting the data to the task space. From the perspective of obtaining random samples, this might not seem like a significant issue, as there are indeed samples close to the nominal solutions. However, the bias introduced in NFs can render the MPPI solution inefficient, which we believe is a key factor contributing to the inferior performance of this method compared to our proposed method, as well as to the MPPI method in certain cases.
@@
(c) Samples from TT-distribution
@@ 6.2 Low-Rank Representation
As highlighted in prior works such as (Shetty et al., 2023), one of the features that enhance the efficiency of TT methods is their ability to represent high-dimensional tensors in low-rank formats. The rank of a tensor can be understood analogously to that of a matrix, defined as the number of independent fibers in the tensor. This low-rank representation plays a pivotal role in our process for several reasons.
First, it allows the entire tensor to be stored in a significantly more memory-efficient manner. Second, techniques such as TT-cross (Savostyanov and Oseledets, 2011) exploit this property to learn the TT representation of a tensor without requiring access to all its elements, reducing computational overhead. Third, this low-rank structure enables faster and more efficient sampling strategies, an advantage leveraged extensively in our proposed method. Finally, it allows operations like tensor products to be executed directly at the core level, bypassing the need to reconstruct or manipulate the full tensor.
@@ 6.3 Limitations
The proposed method has some limitations. The application of TT-SVD may not be well-suited for higher-dimensional systems with fine discretization. TT-SVD assumes the existence of the full tensor before calculating the TT-cores. To address this challenge in our study, we adopted a two-step approach. Initially, a coarse discretization was employed, followed by refinement at the TT-core level using linear or alternative interpolation methods. An alternative strategy to tackle this challenge could involve adopting a multi-resolution discretization approach. For instance, the discretization may be coarse when the robot is distant from objects in the environment and fine when it approaches closer.
In the context of higher-dimensional tasks, an alternative method called TT-cross (Savostyanov and Oseledets, 2011) could be employed instead of TT-SVD. TT-cross aims to learn the cores without constructing the entire tensor. However, for the tasks utilized in this paper, TT-SVD proved sufficient, and we opted for it due to its reliability compared to TT-cross. Future research may explore extensions using the TT-cross method. Another direction could involve combining TT models with other learning algorithms to effectively learn a low-dimensional space, where TT can be employed more efficiently.
@@
Alternatively, recent advances in generative modeling techniques, such as diffusion-based models (Ho et al., 2020) and energy-based models (LeCun et al., 2006), present opportunities for further exploration. These approaches have demonstrated strong performance across various fields, particularly in robotics (Chi et al., 2024; Florence et al., 2021). Recent studies (Gkanatsios et al., 2023; Yang et al., 2023; Du et al., 2020) have also shown their effectiveness within a PoE framework. Furthermore, these methods could be extended for use within the MPC framework discussed in this work, as our PoE-MPPI approach is independent of TT methods and can be adapted to incorporate these generative approaches.
@@
It is also worth noting that while generative methods typically rely on demonstration data to train individual experts, this requirement can be challenging when exploring the entire feasibility region, as demonstrations may not sufficiently cover all possible scenarios. By contrast, the TT approach does not depend on demonstration data and instead operates on functions. This design choice has its advantages for our purpose but may also introduce limitations for certain applications.
Future research should investigate how the TT framework can be combined with these emerging generative models to achieve more robust and scalable solutions. A particularly interesting direction would be exploring how solutions obtained from methods such as MPPI, TT, and diffusion models can be reformulated for seamless integration.
@@
There are two potential avenues for enhancing the proposed method. Firstly, the system’s capability for long-horizon planning could be improved by incorporating an additional expert that considers environmental conditions. These methods can be regarded as supplementary experts and seamlessly integrated with other methodologies through the products of experts method elucidated in this paper, potentially leading to increased efficiency. Notably, recent findings (Shetty et al., 2024; Xue et al., 2024b) demonstrate the applicability of TT models for solving approximate dynamic programming and deriving policy models for robots in TT form. Their approach involves selecting the most optimal or robust action from the model, but we can modify the pipeline by sampling from the TT model and incorporating additional experts, as discussed in this paper. This modification enables a more flexible pipeline, distributing tasks among different experts and mitigating the reliance on achieving convergence to the most optimal solutions.
@@ 7 Conclusion
In this article, we presented a novel approach to enhance the efficiency of sampling-based model predictive control by leveraging products of experts. We illustrated the efficacy of this concept, which differs from traditional projection methods (the sample-then-project strategy or tangent-space-related works) by first projecting the distribution into the feasible area and then sampling from it (the project-then-sample strategy). This approach increases the likelihood of each sample being accepted and provides a more diverse set of samples within the boundaries, rather than focusing solely on the boundaries. Additionally, it is well-suited for tasks where computing the gradient function is challenging, such as non-prehensile manipulation. It also establishes a flexible framework to integrate solutions from different methods, each with distinct objectives and decision variables. This capability enables the construction of a more informative final distribution while simultaneously reducing the complexity of the primary problem.
@@
To address the current limitations of using TT-SVD in higher dimensions, future research could explore alternative TT methods, such as TT-cross, or investigate the potential of initially learning a low-dimensional latent space and subsequently leveraging TT-SVD. Another promising avenue involves incorporating additional experts. For instance, works like (Manchester and Kuindersma, 2020) attempt to directly learn a long-horizon distribution, which could serve as a replacement for the MPPI expert utilized in this article. Similarly, some policies learned in the field of reinforcement learning, such as (Shetty et al., 2024; Xue et al., 2024b), can be viewed as distributions and could function as experts to enhance and modify the proposed method in this study.
{acks}
This work was supported by the SWITCH project (https://switch-project.github.io/), funded by the Swiss National Science Foundation, and by the China Scholarship Council (grant No.202106230104).
@@ A.1 TT-Sample
A.2 Proj-MPPI
@@ A.3 NFs-PoE-MPPI
Param. Obstacle Avoidance Push. Cyl. Push. Box Push. Mus. Sphere man. Sin. man.
##\## data $4\times 10^{6}$ $10^{6}$ $10^{7}$ $10^{7}$ $\approx 1.5\times 10^{7}$ $\approx 1.5\times 10^{7}$
TT ##\## State Discretization $100$ $10$ $10$ $10$ $25$ $25$
##\## Action Discretization $20$ $10$ $10$ $10$ $10$ $10$
State Refinement Scale $1$ $20$ $10$ $10$ $10$ $10$
Action Refinement Scale $10$ $20$ $10$ $10$ $10$ $10$
NFs ##\## Coupling Layers $6$ $8$ $8$ $8$ $8$ $6$
Size of layers $256$ $256$ $256$ $256$ $256$ $256$
Batch size $4096$ $4096$ $8192$ $8192$ $8192$ $8192$
learning rate $0.0004$ $0.0002$ $0.0001$ $0.0001$ $0.0002$ $0.0004$
A.4 Experiments

### v2

@@ 6 Discussion
The primary objective of this article is to enhance the sampling distribution employed in SMPC through the application of products of experts in the solution space. As anticipated, all methods with TT-PoE pipeline (NFs-PoE-MPPI, TT-PoE-MPPI, and TT-PoE-DIAL) exhibit superior performance compared to their conventional counterparts. However, the performance of NFs-PoE was not significantly better than that of Proj-MPPI, and in all experiments, TT-PoE consistently outperformed NFs-PoE by a substantial margin. In the subsequent sections, we will delve deeper into the reasons behind this performance, explaining the advantages of using TT-distribution over NFs models and identifying its limiting factors.
@@ 6.1 TT-distribution vs Normalizing flow
Nevertheless, similar to other latent space learning methods, the acquired latent space lacks intuitiveness (Losey et al., 2022). In this context, intuitive refers to the observed relation between data points in the latent space, not necessarily aligning with their distance in the main spaces. For instance, when two proximate data points exhibit similar distances to an objective point in latent space, while in task space, one is positioned near the nominal solution and the other is distant. This phenomenon is illustrated in Fig. 11, obtained from a model trained for the PNGRID task, where the nominal action is presumed to be [0.5,0.5][-0.5,-0.5] at state [1,0.9][-1,-0.9], and the Gaussian model utilized for the MPPI method possesses identical variance in both directions. At this point, only velocity constraints (ranging between -1 and 1) must be active, with no other constraints in effect.
@@
Observing the task space samples, it is evident that they are not entirely clustered around the nominal action [0.5,0.5][-0.5,0.5]. Additionally, the distribution of samples in both directions appears dissimilar. This discrepancy arises due to the nonlinear function employed to revert the data back to the task space, which may have varying effects in different directions. This suggests that the covariance matrix in the latent space must be meticulously chosen to ensure the desired distribution upon reverting the data to the task space. From the perspective of obtaining random samples, this might not seem like a significant issue, as there are indeed samples close to the nominal solutions. However, the bias introduced in NFs can render the MPPI solution inefficient, which we believe is a key factor contributing to the inferior performance of this method compared to our proposed method, as well as to the MPPI method in certain cases.
@@
(c) Samples from $T$-distribution
@@ 6.2 Low-Rank Representation
As highlighted in prior works such as (Shetty et al., 2024a), one of the features that enhance the efficiency of TT methods is their ability to represent high-dimensional tensors in low-rank formats. The rank of a tensor can be understood analogously to that of a matrix, defined as the number of independent fibers in the tensor. This low-rank representation plays a pivotal role in our process for several reasons.
First, it allows the entire tensor to be stored in a significantly more memory-efficient manner. Second, techniques such as TT-cross (Savostyanov and Oseledets, 2011) exploit this property to learn the TT representation of a tensor without requiring access to all its elements, reducing computational overhead. Third, this low-rank structure enables faster and more efficient sampling strategies (compared to sampling directly from the whole tensor), an advantage leveraged extensively in our proposed method. Finally, it allows operations like tensor products to be executed directly at the core level, bypassing the need to reconstruct or manipulate the full tensor.
@@
6.3 Computation time
As outlined in Alg. 1, our method adds only two extra operations (lines 3–4) and modifies the sampling step (line 5) compared to MPPI. Table 6 shows how these changes affect computation time. Although MPPI achieves faster overall runtimes, our TT-sampling times are on par with other baselines, making it suitable for online planning.
In its current Python implementation, TT‑sampling scales linearly with the number of action variables—e.g., the whole‑body obstacle‑avoidance task uses 3.5× more—because it relies on a Python library (i.e., tntorch) with a for‑loop inside, and lacks multithreading. By contrast, MPPI can leverage highly optimized native libraries and its state‑independent Gaussian distribution to improve the computation time significantly. By migrating performance‑critical TT sections to C/C++ and enabling multithreading, we estimate TT’s sampling time could drop below 1 ms. Until those optimizations are implemented, the timings reported here reflect the current implementation. All measurements were recorded using an NVIDIA RTX 3090 GPU and a 12‑core CPU.
Table 6: Average computation times (in milliseconds) for each method on the PNGRID (P) and whole-body obstacle avoidance (W) tasks. MPPI’s runtimes were virtually identical across both tasks and are therefore reported once. Note that computation times for related methods (e.g., DIAL and TT-PoE-DIAL) can be inferred directly from this table. Although our proposed method is slower than MPPI, it is on par with the other techniques and, as shown in the performance tables, yields a significant improvement over MPPI.
Method Product [ms] Conditioning [ms] Sampling [ms] Mapping [ms] Projection [ms] Total [ms]
NFs-PoE-MPPI (P) $0.03\pm 0.00$ - $0.48\pm 0.02$ $2.50\pm 0.07(\times 2)$ - $5.51\pm 0.16$
MPPI - - $0.48\pm 0.02$ - - $\bm{0.48\pm 0.02}$
Proj_MPPI (P) - - $0.48\pm 0.02$ - $02\pm 26.46 5.5$$50\pm 26.48$
TT-PoE-MPPI (P) $02\pm 0.00 0.0$$04\pm 0.00$ $2.67\pm 0.02$ - - $2.73\pm 0.02$
TT-PoE-MPPI (W) $0.13\pm 0.01$ $1.23\pm 0.10$ $5.92\pm 0.31$ - - $7.28\pm 0.42$
@@ 6.4 Limitations
Despite its strengths, our approach has scalability limits: while we’ve demonstrated performance on moderately high-dimensional systems (which can be sufficient for a wide range of robotic tasks), it does not extend readily to very high dimensions (e.g., over 50). Several avenues could address this. First, a multi-resolution discretization scheme (e.g., using coarse grids when the robot is far from obstacles and finer grids when it nears them) can dramatically shrink the effective discretization size and enable higher-dimensional problems. Second, leveraging cumulative learning to inform the TT-cross algorithm may accelerate convergence and improve accuracy. Finally, integrating TT models with complementary learning techniques to discover compact latent representations could further reduce dimensionality and allow TT methods to operate more efficiently in even larger state spaces.
@@
Alternatively, recent advances in generative modeling techniques, such as diffusion-based models (Ho et al., 2020) and energy-based models (LeCun et al., 2006), present opportunities for further exploration. These methods could be extended for use within the MPC framework discussed in this work, as our PoE-MPPI approach is independent of TT methods and can be adapted to incorporate these generative approaches. In this work, we have already demonstrated the mutual benefits of combining our pipeline with DIAL-MPC (a data-free, diffusion-inspired sampler), and future work could similarly integrate other generative approaches to further improve sampling efficiency and constraint satisfaction.
@@
It is also worth noting that while generative methods typically rely on demonstration data to train individual experts, this requirement can be challenging when exploring the entire feasibility region, as demonstrations may not sufficiently cover all possible scenarios. By contrast, the TT approach does not depend on demonstration data and instead operates on functions. This design choice has its advantages for our purpose but may also introduce limitations for certain applications. In some applications, a hybrid strategy that combines data-driven and function-based experts could offer the best of both worlds.
@@
There are two potential avenues for enhancing the proposed method. Firstly, the system’s capability for long-horizon planning could be improved by incorporating an additional expert that considers environmental contexts. These methods can be regarded as supplementary experts and seamlessly integrated with other methodologies through the products of experts method elucidated in this paper, potentially leading to increased efficiency. Secondly, recent findings (Shetty et al., 2024b; Xue et al., 2024b) demonstrate the applicability of TT models for solving approximate dynamic programming and deriving policy models for robots in TT form. Their approach involves selecting the most optimal or robust action from the model, but we can modify the pipeline by sampling from the TT model and incorporating additional experts, as discussed in this paper. This modification enables a more flexible pipeline, distributing tasks among different experts and mitigating the reliance on achieving convergence to the most optimal solutions.
@@ 7 Conclusion
In this article, we presented a novel approach to enhance the efficiency of sampling-based model predictive control by leveraging products of experts in the solution space and discussing the introduction of a feasibility expert. We illustrated the efficacy of this concept, which differs from traditional projection methods (the sample-then-project strategy or tangent-space-related works) by first projecting the optimality distribution into the feasible area and then sampling from it (the project-then-sample strategy). This approach increases the likelihood of each sample being accepted and provides a more diverse set of samples within the boundaries, rather than focusing solely on the boundaries. Additionally, it is well-suited for tasks where computing the gradient function is challenging, such as non-prehensile manipulation. It also establishes a flexible framework to integrate solutions from different methods, each with distinct objectives and decision variables. This capability allows us to reshape the sampling distribution before drawing samples, enabling more targeted and efficient exploration. We demonstrated how this advantage can be leveraged across different sampling-based frameworks, including MPPI and DIAL-MPC, resulting in improved sampling efficiency and task performance in both cases.
@@
To address the current limitations of our method in higher dimensions, future research could investigate the potential of initially learning a low-dimensional latent space and subsequently leveraging the TT-PoE method. Another promising avenue involves incorporating additional experts with additional modalities (e.g., demonstrations or other optimality distributions . Similarly, some policies learned in the field of reinforcement learning, such as (Shetty et al., 2024b; Xue et al., 2024b), can be viewed as distributions and could function as experts to enhance and modify the proposed method in this study.
Declaration of Conflicting Interests
The authors declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
Funding
The authors disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: This work was supported by the State Secretariat for Education, Research and Innovation in Switzerland for participation in the European Commission’s Horizon Europe Program through the INTELLIMAN project (https://intelliman-project.eu/, HORIZON-CL4-Digital-Emerging Grant 101070136) and the SESTOSENSO project (https://sestosenso.eu/, HORIZON-CL4-Digital-Emerging Grant 101070310), as well as by the China Scholarship Council (grant No. 202106230104).
ORCID iDs
@@ A.1 TT-Sample
A.2 Conditional TT
Let assume that $\bm{x}$ be composed of two parts as $\bm{x}=(\bm{y}_{1},\bm{y}_{2})\in\Omega_{\bm{x}}$, where $\bm{y}_{1}$ and $\bm{y}_{2}$ refers to the first $d_{1}$ and the last $d-d_{1}$ variables in $\bm{y}$, respectively. Our goal is to find $\text{Pr}(\bm{y}_{2}|\bm{y}_{1})$ of the TT distribution given in (8).
For a fixed value $\bm{y}_{1}=\bm{y}_{t}$, we define a new TT model $\bm{\mathcal{P}}^{\bm{y}_{t}}$ over $\bm{y}_{2}$ by slicing the original cores:
$\bm{\mathcal{P}}^{\bm{y}_{t}}_{\bm{y}_{2}}\;=\;\bm{\mathcal{P}}_{(\bm{y}_{t},\bm{y}_{2})}\quad\forall\;\bm{y}_{2}\in\Omega_{\bm{y}_{2}}.$
Equivalently, its cores become
$(\bm{\mathcal{P}}^{\bm{y}_{t}})^{k}\;=\;\begin{cases}\bm{\mathcal{P}}^{k}_{:,\,y_{t_{k}},\,:},&k=1,\dots,d_{1},\\[6.0pt] \bm{\mathcal{P}}^{k},&k=d_{1}+1,\dots,d.\end{cases}$ (27)
The conditional probability then follows as
$\Pr(\bm{y}_{2}\mid\bm{y}_{1}=\bm{y}_{t})\;=\;\frac{\bigl|\bm{\mathcal{P}}_{\bm{y}_{2}}^{\bm{y}_{t}}\bigr|}{Z_{1}}\quad\forall\;\bm{y}_{2}\in\Omega_{\bm{y}_{2}}.$ (28)
Finally, one can draw samples of $\bm{y}_{2}$ given $\bm{y}_{1}=\bm{y}_{t}$.
A.3 Proj-MPPI
@@ A.4 NFs-PoE-MPPI
Param. Obstacle Avoidance Push. Cyl. Push. Box Push. Mus. Sphere man. Sin. man. Whole-body
$T$ #\# State Discretization $100$ $10$ $10$ $10$ $25$ $25$ $50$
#\# Action Discretization 2020 $10$ $10$ $10$ $10$ $10$ $50$
State Refinement Scale 11 2020 $10$ $10$ $10$ $10$ 11
Action Refinement Scale $10$ 2020 $10$ $10$ $10$ $10$ 11
NFs #\# Coupling Layers 66 88 88 88 88 66 -
Size of layers $256$ $256$ $256$ $256$ $256$ $256$ -
Batch size $4096$ $4096$ $8192$ $8192$ $8192$ $8192$ -
learning rate $0.0004$ $0.0002$ $0.0001$ $0.0001$ $0.0002$ $0.0004$ -
A.5 MPPI on the Sphere Manifold
For the intrinsic method used in Section 5.2, we adopt the exponential map to enable sampling in the tangent space of the sphere, followed by mapping the samples back to the manifold. At each timestep, the tangent space is defined at a point $\bm{x}_{t}\in\mathcal{M}=\left\{\bm{x}\in\mathbb{R}^{3}\,\middle|\,\|\bm{x}\|=r\right\}$ on the sphere. Any valid action $\bm{u}_{t}\in\mathbb{R}^{3}$ must lie in this tangent space, and thus satisfy the orthogonality condition $\bm{x}_{t}^{\top}\bm{u}_{t}=0$.
This tangent space is a 2D subspace of $\mathbb{R}^{3}$, and we generate actions $\bm{u}_{t}$ by first sampling in this lower-dimensional space: $\bm{u}_{t}^{2D}\sim\mathcal{N}(\bm{\mu}_{t},\bm{\Sigma}_{t})$. The mapping from the 2D tangent coordinates to the 3D ambient space is defined as
$\bm{u}_{t}=\bm{T}(\bm{x}_{t})\bm{u}_{t}^{2D},$ (30)
where the transformation matrix $\bm{T}(\bm{x}_{t})=[\bm{e}^{\top}_{1}(\bm{x}_{t}),\bm{e}^{\top}_{2}(\bm{x}_{t})]^{\top}$ is constructed from an orthonormal basis of the tangent space at $\bm{x}_{t}$
$\bm{e}_{1}=\frac{\bm{l}\times\bm{x}_{t}}{\|\bm{l}\times\bm{x}_{t}\|},\quad\bm{e}_{2}=\frac{\bm{e}_{1}\times\bm{x}_{t}}{\|\bm{e}_{1}\times\bm{x}_{t}\|}.$ (31)
Here, $\bm{l}$ is an arbitrary vector chosen to avoid alignment with $\bm{x}_{t}$, ensuring a well-defined cross product
$\bm{l}=\begin{cases}[1,0,0]^{\top}&\text{if }|\bm{x}_{t}^{\top}[1,0,0]^{\top}|\leq 0.9,\\ [0,1,0]^{\top}&\text{otherwise}.\end{cases}$ (32)
After sampling actions and mapping them to the task space, the next state $\bm{x}_{t+1}$ is computed using the exponential map on the sphere
$\bm{x}_{t+1}=\exp_{\bm{x}_{t}}(\bm{v})=\cos\left(\frac{\|\bm{v}\|}{r}\right)\bm{x}_{t}+\sin\left(\frac{\|\bm{v}\|}{r}\right)\frac{r\,\bm{v}}{\|\bm{v}\|},$ (33)
where $\bm{v}=\bm{u}_{t}\cdot dt$ is the scaled action vector, and $r$ is the radius of the sphere. This formulation ensures that the resulting state $\bm{x}_{t+1}$ remains on the sphere manifold.
A.6 DIAL-mpc and TT-PoE-DIAL
Inspired by the success of diffusion-based methods across various domains, Pan et al. (2024) proposed a novel MPPI variant for trajectory optimization. In their approach, each MPPI iteration is repeated $N_{\text{diffuse}}$ times using different noise realizations. Their results (which we also observe in our own experiments) demonstrate that this method significantly outperforms vanilla MPPI and several other baselines. Building on this idea, Xue et al. (2025) further extended the method to a full MPC framework.
In their formulation, the noise covariance at each diffusion step is defined as
$\bm{\Sigma}^{n}_{t+h}=\exp\left(-\frac{N_{\text{diffuse}}-n}{\beta_{1}N_{\text{diffuse}}}-\frac{H-h}{\beta_{2}H}\right)\bm{I},$ (34)
where $\beta_{1}$ and $\beta_{2}$ are temperature parameters controlling the diffusion process and the time horizon, respectively. For simplicity, we define two decay factors: $f_{\text{denoise}}=\exp\left(-\frac{1}{\beta_{1}N_{\text{diffuse}}}\right)$ and $f_{\text{horizon}}=\exp\left(-\frac{1}{\beta_{2}H}\right)$.
Algorithm 2 outlines how this diffusion-based MPPI method can be integrated into the TT-PoE pipeline. The only modifications required are the addition of lines 5 and 6, and replacing Gaussian sampling with the TT-based sampling approach described in Equation (20).
Algorithm 2 $T$-PoE-DIAL
1:Require: $\bm{x}_{t},\bm{\mu}_{t,\ldots,t+H-1},\bm{\sigma}_{t,\ldots,t+H-1}$
2: Return: $\bm{u}^{*}_{t}$
[…]
4:for $n$ from 0 to $N_{\text{diffuse}}$ do
5:  for $h$ from 0 to H1H-1 do
6:   Update the covariance matrix $\triangleright$ Eq. (34)
7:   Calculate $\bm{\mathcal{P}}_{\mathcal{N}_{m}}$ $\triangleright$ Eq. (21)
8:   $\bm{\mathcal{P}}_{h}\leftarrowProduct$ of the distributions $\triangleright$ Eq. (20)
9:   $\hat{\bm{u}}^{0}_{t+h},\ldots,\hat{\bm{u}}^{N}_{t+h}\leftarrowSample$ NN actions
10:   $\hat{\bm{x}}^{0}_{t+h+1},\ldots,\hat{\bm{x}}^{N}_{t+h+1}\leftarrowForward$ Dynamics
11:  end for
12:  $c^{0},\dots,c^{N}\leftarrowCalculate$ cost values
13:  $\bm{\mu}_{t}\leftarrow$ update the mean value $\triangleright$ Eq. (4)
14:end for
[…]
A.7 Experiments